Two points P and Q are maintained at the potentials of 10V and –4V, respectively. The work done in moving 100 electrons from P to Q is ?

Grade:11

2 Answers

Arun
3 years ago
potential difference=V=410=14V
 
Charge=q=100e=1.6×1017C
 
Now, potential difference, V=qW
 
W=qV=1.6×1017×(14)
 
W=2.24×1016J
Vikas TU
3 years ago
We know , work done in moving a charge q is given by :
W=q(Vf-Vi)     ......( 1 )
Here :
q = 100e 
Putting all value in equation 1 :
We get :
W=-1.6* 10^{-17}*(-4-10)
W=1.6*14*10^{-17} J
W=2.24 * 10^{-16} J
Therefore , this is the required solution .